# KNighter: Transforming Static Analysis with LLM-Synthesized Checkers

### v1

@@ 5.3 RQ3: Baseline Comparison
Since no comparable automated static analyzer generation approaches exist for this domain, we evaluate KNighter against expert-written checkers. Our baseline is provided by Smatch [1], which is widely used in Linux kernel analysis and supports tailored checks for all bug types considered in Table 1. We conducted the comparative analyses by running Smatch on the entire codebase to determine if it could detect the bugs found by KNighter.
@@
Our evaluation revealed that Smatch reported 1970 errors and 2870 warnings across the kernel. Notably, it failed to detect any of our true positive bugs, underscoring KNighter’s unique detection capabilities. Further investigation of Smatch’s checkers showed they don’t fully leverage domain-specific knowledge available in the Linux kernel—knowledge that KNighter extracts from historical patches. For instance, Smatch’s check_deref checker employs static range analysis to identify potential null pointers but lacks domain-specific insights. It fails to recognize that functions like devm_kzalloc may return NULL under error conditions that conventional static range analysis cannot detect. Consequently, Smatch identified only three potential null pointer dereferences, all of which were confined to unit test files rarely prioritized by developers. We conclude that KNighter and Smatch detect different classes of bugs, demonstrating KNighter’s effectiveness in learning domain knowledge from patches and subsequently identifying diverse bugs and vulnerabilities.
5.4 RQ4: Effectiveness of Components
5.4.1 Bug Triage Agent
To evaluate our triage agent, from the valid checkers, we sampled up to 5 reports per checker to reduce manual inspection efforts. In total, we collected 79 reports from 18 checkers. The remaining 21 valid checkers did not identify any bugs, while 2 non-plausible checkers also generated reports. These reports may include those generated during the refinement process. Our objective here was to evaluate the effectiveness of our triage agent. Among these reports, 29 were labeled as “bug” (positive) by our triage agent, while the remaining 50 were labeled as “not-a-bug” (negative). Two of our authors manually cross-checked all reports to establish ground truth. Our triage agent achieved 7 true positives (TP), 22 false positives (FP), 50 true negatives (TN), and zero false negatives (FN). Notably, the absence of false negatives is beneficial because it helps ensure that most true bugs are captured, even if this comes with some false positives that can be filtered out in subsequent steps or human inspection.
5.4.2 Ablation Study
@@
Table 3: Ablation study results. “Default” means KNighter’s standard configuration utilizing multi-stage synthesis and the O3-mini language model. Alternative configurations are compared against this baseline.
@@
To further evaluate our design choices, we created a sample dataset of patch commits for an ablation study. We randomly sampled 2 commits from each bug type using zero as the random seed. This resulted in a dataset of 20 commits (2 commits × 10 bug types). Table 3 shows the results of our ablation study.
@@
Checker synthesis. Our default pipeline implements a three-stage approach: bug pattern analysis, detection plan synthesis, and checker implementation. We evaluated the effectiveness of directly synthesizing checkers without the initial bug pattern analysis and plan synthesis phases while maintaining identical in-context learning examples across configurations. As shown in Table 3, without multi-stage synthesis, KNighter produced valid checkers for only 8 commits compared to 12 with our default approach. The single-stage approach also generated significantly more syntax errors (52 vs. 28), resulting in checkers that failed to compile.
@@
LLM choice. We evaluated KNighter’s performance across different language models for checker synthesis. In addition to our default model O3-mini (reasoning model), we tested GPT-4o and Gemini-2-flash. As Table 3 demonstrates, GPT-4o performed slightly worse than O3-mini, generating 11 valid checkers compared to 12. Gemini-2-flash performed substantially worse, producing valid checkers for only 4 commits. Upon closer inspection, we found that Gemini-2-flash struggled with Clang Static Analyzer implementation, frequently using non-existent APIs and generating syntax errors at a much higher rate (130 vs. 28).
6 Related Work
6.1 Traditional Static Analysis
@@
Given the cruciality of Linux kernel and the diversity of its vulnerabilities, many static analyzers have been developed to target different classes of bugs.
CRIX [20] focuses on detecting missing-check bugs in kernel by constructing def-use slices of critical variables. Goshawk [21] proposes a structure-aware and object-centric abstraction over memory management operations to aid detecting double-free and use-after-free bugs. UBITect [39] integrates type qualifier analysis and symbolic execution to detect use-before-initialization bugs in the Linux kernel, which might encounter challenges due to exhaustive path exploration. CRed [34] detects Use-After-Free bugs in scale by spatio-temporal context reduction. LR-Miner [16] establishes field-sensitive locking rules between locks and variables. Then it detects rule violations to pinpoint data races in mainstream OS kernels. Authors of [12] extensively analyze over one thousand reference counting bugs in the history of Linux kernel and derived 9 anti-patterns. They further developed a set of static analyzers to detect such patterns. LinkRID [19] also targets refcounting bugs by modelling common Linux-specific refcount usage conventions. DCUAF[5] conducts local-global analysis to find concurrent functions to discover kernel currency bugs. SUTURE[41] performs high-order taint analysis to discover kernel vulnerabilities caused by malicious userspace input.
While traditionally designed static analyzers can perform well on their targeted classes of bugs, they require extensive analysis and implementation effort from human experts. This limits their ability to scale to additional bug classes and keep pace with the rapid evolution of the Linux kernel. KNighter is complementary to such analyzers by detecting more classes of bugs and adapting to newly developed bug patterns.
Beyond one specific class of bugs, many static analysis systems implement different analysis techniques to improve precision and efficiency. FiTx [28] implements fast analysis for single-compilation unit without the requirement of complex features like indirect call analysis. PATA [17] improves the alias analysis precision using control-flow paths and access paths information. FiTx and PATA support detecting 6 and 3 common bug patterns respectively, by encoding bug patterns as finite state machines. While these static analysis systems can potentially support more classes of bugs, analyzing bugs and designing the pattern state machines remain challenging and labor-intensive. KNighter automates this process by automatically synthesizing bug patterns from patches. Beyond automation, although KNighter generates CSA checkers, the idea is generalizable to using other static analysis tools without fundamental obstacles since LLMs have been trained on most existing tools.
6.2 LLM-Based Static Analysis
With the recent advances of LLMs, many techniques leverage LLMs’ code comprehension ability to mediate the limitations of traditional static analysis tools that typically require approximation or expert-produced knowledge. IRIS [18] enhanced CodeQL’s taint analysis by using LLMs to infer taint specifications for external APIs, which would have required manual labeling. To detect Use-Before-Initialization bugs, LLift [15] uses LLMs to extract initializers and generate their post-constraints. The post-constraints are then used during further analysis to prune unreachable paths. InferROI [30] uses LLMs to infer resource-oriented intentions and then perform lightweight static resource leak detection. While the above approaches can enhance static analyzers, a substantial portion of the analyzer still needs to be crafted by researchers. In contrast, KNighter can automatically synthesize the entire checker without human intervention. Beyond this, the checkers generated by KNighter could potentially be further complemented by additional information provided by the above approaches.
Another branch of work directly use LLMs to analyze the code without traditional static analysis tools. Vul-RAG [10] constructs a knowledge base from existing CVE instances and directly queries LLMs with Retrieval-Augmented-Generation to determine whether a new piece of code has the same bug. Zhang et al. [40] studies various prompt design techniques for vulnerability detection. Fine-tuning has also been explored by [25] to perform vulnerability detection. The above LLM-Based analysis techniques require querying LLMs with the code under test. In the context of Linux kernel, querying LLMs with millions lines of code, potentially multiplied by the number of different classes of bugs, is prohibitively expensive. In comparison, KNighter generates a static analysis checker with negligible token cost that can be used to efficiently scan the kernel codebase.
7 Conclusion
@@
This paper introduces KNighter, a novel approach that transforms how LLMs can contribute to static analysis for complex systems like the Linux kernel. By synthesizing specialized static analyzers rather than directly analyzing code, KNighter bridges the gap between LLMs’ reasoning capabilities and the practical constraints of analyzing massive codebases. KNighter’s practical impact is demonstrated by the discovery of 70 previously unknown bugs in the Linux kernel, with 56 confirmed and 41 already fixed.
@@
Looking forward, KNighter opens new possibilities for practical LLM-based static analysis. Future work could extend this approach to other systems beyond the Linux kernel, incorporate additional learning paradigms, and further refine checker generation techniques to address more complex bug patterns. By leveraging LLMs to synthesize tools rather than perform analysis directly, we establish a scalable, reliable, and traceable paradigm for utilizing AI in critical software security applications.

### v2

@@ 5.3. RQ3: Orthogonality with Smatch
Since no comparable automated static analyzer generation approaches exist for this domain, we evaluate KNighter against expert-written checkers. Our baseline is provided by Smatch (smatch, ), which is widely used in Linux kernel analysis and supports tailored checks for all bug types considered in Table 1. We conducted the comparative analyses by running Smatch on the entire codebase to determine if it could detect the bugs found by KNighter.
@@
Smatch reported a total of 1970 errors and 2870 warnings across the kernel. We manually inspected all the files where bugs were detected to assess whether any of the true positive bugs identified by KNighter were also detected by Smatch. Notably, Smatch failed to detect any of our true positive bugs, underscoring the unique detection capabilities of KNighter. Further analysis of Smatch’s checkers revealed that they do not fully leverage the domain-specific knowledge embedded in the Linux kernels—a resource that KNighter effectively extracts from historical patches. For instance, Smatch’s check_deref checker employs static range analysis to identify potential null pointers but lacks domain-specific insights. It fails to recognize that functions like devm_kzalloc may return NULL under error conditions that conventional static range analysis cannot detect. Consequently, Smatch identified only three potential null pointer dereferences, all of which were confined to unit test files rarely prioritized by developers. We conclude that KNighter and Smatch detect different classes of bugs, demonstrating KNighter’s effectiveness in learning domain knowledge from patches and subsequently identifying diverse bugs and vulnerabilities.
5.4. RQ4: Effectiveness of Components
5.4.1. Bug Triage Agent
To evaluate our bug triage agent, from the 39 valid checkers, we sampled up to 5 reports per checker to reduce manual inspection efforts, aiming to reduce manual inspection efforts while maintaining evaluation coverage. In total, we collected 79 reports from 18 checkers, while the remaining 21 valid checkers didn’t generate reports. Our triage agent classified 29 reports as “bug” (positive) and 50 as “not-a-bug” (negative). These classifications were compared against ground truth labels established by manual review from two authors. The agent achieved 7 true positives (TP), 22 false positives (FP), 50 true negatives (TN), and zero false negatives (FN). The absence of false negatives is particularly important, as it indicates the agent effectively prioritized all potentially true bugs in this set, minimizing the risk of overlooking genuine issues, even though the 22 false positives require further filtering.
5.4.2. Ablation Study
@@
Table 2. Ablation study results. “Default” means KNighter’s standard configuration utilizing multi-stage synthesis, fixed few-shot examples, and the O3-mini model. Alternative configurations are compared against this baseline.
@@
To evaluate our design choices, we created a sample dataset of patch commits for an ablation study. We randomly sampled 2 commits from each bug type using zero as the random seed. This resulted in a dataset of 20 commits (2 commits × 10 bug types). Table 2 shows the overall results of our ablation study.
@@
Checker synthesis. First, we assessed the impact of our default three-stage synthesis approach (bug pattern analysis, plan synthesis, checker implementation) compared to directly synthesizing checkers in a single stage (omitting explicit pattern/plan steps), using identical few-shot examples. As shown in Table 2, the multi-stage approach proved more effective, yielding valid checkers for 12 commits compared to only 8 for the single-stage method. Furthermore, the single-stage approach resulted in significantly more syntax errors (52 vs. 28), often leading to checkers that failed to compile. This highlights the value of the structured multi-stage process for improving both validity and compilability.
Second, we explored using Retrieval-Augmented Generation (RAG) (gao2023retrieval, ) for selecting few-shot examples dynamically, comparing it against our default set of fixed examples. It utilizes a knowledge base derived from 118 official CSA checkers (clang-static-analysis, ), embedded using text-embedding-ada-002 (text-embedding, ). During synthesis, three relevant examples are retrieved based on semantic similarity. Our results indicate that RAG-based example selection achieved comparable effectiveness to our fixed examples, also generating valid checkers for 12 commits. However, because the official CSA checkers used are substantially longer than our curated fixed examples, the RAG approach incurred approximately double the input/output token cost. Due to similar effectiveness, our default fixed few-shot examples offer better cost-efficiency for this task.
LLM choice. We evaluated the performance of KNighter across different language models for checker synthesis. In addition to our default model, O3-mini, we tested GPT-4o, Gemini-2-flash, and open-source DeepSeek-R1.
As detailed in Table 2, O3-mini yielded the most valid checkers (12). GPT-4o and DeepSeek-R1 performed comparably to each other, generating 11 valid checkers each, only slightly fewer than O3-mini. This suggests that multiple high-capability models, including open-source options, are viable, although minor differences in performance exist.
@@
In contrast, Gemini-2-flash performed substantially worse, producing valid checkers for only 4 commits. Upon closer inspection, we found that Gemini-2-flash struggled with CSA implementation, frequently using non-existent APIs and generating syntax errors at a much higher rate (130 vs. 28). This highlights a crucial insight: successful checker synthesis demands more than general coding proficiency; accurate knowledge or inference of the target framework’s specific APIs and conventions (CSA in this case) is essential.
6. Related Work
6.1. Traditional Static Analysis
@@
Given the cruciality of the Linux kernel and the diversity of its bugs, many static analyzers have been developed. These generally fall into several categories based on their approach.
Rule/Model-based analyzers. A significant body of work focuses on detecting specific classes of bugs using predefined rules or models. Examples include CRIX (kernel-static-lu2019detecting, ) (detecting missing checks via def-use slices), Goshawk (kernel-static-lyu2022goshawk, ) (memory corruption analysis), UBITect (zhai2020ubitect, ) (Use Before Initialization bugs), CRed (yan18uaf, ) (Use-After-Free detection), LR-Miner (li2024lr, ) (data races via locking rules), tools targeting refcounting bugs based on derived anti-patterns (kernel-static-refcounting, ) or specific conventions (liu22linkrid, ), DCUAF (bai19uaf, ) (concurrent Use-After-Free), and SUTURE (zhang21suture, ) (taint analysis for userspace input vulnerabilities). While often effective for their targeted bug classes, these approaches typically require extensive human expertise for analysis design and implementation, limiting their scalability and adaptability to newly emerging bug patterns in the rapidly evolving kernel.
Patch-based specification inference. Another line of work leverages historical patches to infer specifications, which are then used by separate checkers. For instance, APHP (aphp, ) extracts API Post-Handling (APH) specifications from both code and descriptions in patches to detect APH violations. A very recent work, Seal (seal, ), analyzes security patches to infer diverse specifications for Linux interfaces. While powerful, these methods generally rely on existing static analysis infrastructure to enforce the inferred specifications. We attempted to include Seal in our experimental comparison, but encountered practical difficulties as its publicly available version depends on private commercial tools and was not compatible with the recent Linux kernel versions.
General static analysis frameworks. Other systems focus on improving underlying static analysis techniques for better precision or efficiency across multiple bug types. FiTx (suzuki2024fitx, ), for example, implements fast analysis for single compilation units, while PATA (li22pata, ) enhances alias analysis precision using path information. These systems often encode bug patterns as state machines (suzuki2024fitx, ; li22pata, ), but designing these patterns still typically involves manual effort.
Our synthesis approach. In contrast to the above, our work, KNighter, uses LLMs to synthesize the static analyzer itself directly from historical patch information. Instead of relying on predefined rules or inferring specifications for existing checkers, KNighter learns both the bug pattern and the corresponding detection logic, generating a new checker automatically. This approach fundamentally differs by automating the analyzer creation process, potentially offering greater adaptability to new bug types compared to traditional methods or specification inference. This distinction may also make KNighter more versatile; while specification-inference tools excel at finding bugs related to the interfaces they model, KNighter can synthesize checkers for a broader range of issues demonstrated in patches, including problems not strictly tied to interface misuse, such as certain integer overflows or complex logic errors. Furthermore, although KNighter currently generates CSA checkers, the core idea of LLM-based checker synthesis is potentially generalizable to target other static analysis frameworks, given the broad training data of modern LLMs. Thus, KNighter complements existing approaches by providing a mechanism to automatically generate analyzers for newly observed bug patterns found in patches.
6.2. LLM-Based Static Analysis
Recent LLM advancements enable new static analysis techniques, differing significantly from our synthesis approach.
LLM-augmented static analysis. Some techniques use LLMs to assist existing tools by automating tasks previously requiring manual effort. For example, LLMs can infer taint specifications for external APIs (IRIS (li2024llm, )), generate post-constraints to prune analysis paths (LLift (li2024enhancing, )), or infer resource-handling intentions (InferROI (wang2024resourceleak, )). While helpful, these methods still fundamentally depend on a substantial, human-developed analyzer core. This can limit their ability to easily adapt or generalize to detect novel bug types beyond those originally targeted by the core analyzer’s design.
Direct code analysis with LLMs. Other work uses LLMs to directly analyze source code, using techniques like Retrieval-Augmented Generation (Vul-RAG (du2024vul, )), prompt engineering (cy24prompt, ), or fine-tuning (shestov, ). However, directly applying LLMs to scan large systems like the Linux kernel is often prohibitively expensive and faces scalability challenges.
Our synthesis approach. In contrast, KNighter uses LLMs to synthesize the entire static analyzer from patches. Unlike LLM-augmented tools, this minimizes reliance on pre-existing manual analyzer development, enhancing adaptability for diverse bug types. Unlike direct LLM analysis, KNighter generates efficient, reusable static checkers, avoiding the high costs and scalability issues of scanning massive systems directly with LLMs. This makes KNighter a practical approach for applying LLM intelligence to large-scale static analysis, complementing other methods by automatically creating analyzers for new bug patterns. Checkers generated by KNighter could also potentially benefit from integration with LLM-augmented techniques.
7. Conclusion
@@
This paper introduces KNighter, a novel approach that transforms how LLMs can contribute to static analysis for complex systems like the Linux kernel. By synthesizing static analyzers rather than directly analyzing code, KNighter bridges the gap between LLMs’ reasoning capabilities and the practical constraints of analyzing massive systems. KNighter’s practical impact is shown by the discovery of 92 new, long-latent bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 16 CVE assigned.
@@
Looking forward, KNighter opens new possibilities for scalable LLM-based static analysis. Future work could extend this approach to other systems beyond the Linux kernel, incorporate additional learning paradigms, and further refine checker generation techniques to address more complex bug patterns. By leveraging LLMs to synthesize tools rather than perform analysis directly, we establish a scalable, reliable, and traceable paradigm for utilizing AI in critical software security applications.